# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ 4.4 VMS feedback tests
Therefore, the only remaining strong feedback mechanism has to do with radiation. In the third test, we disable the radiative feedback from all stars (the VMS population alone elevates the overall L/M ratio, but only by a factor of $\sim 1.5$, see § 3). With the third row in Fig. 12, we see that radiative feedback indeed makes a big difference. The SFE is much higher as the cumulative stellar mass reaches $\sim 2.5\times 10^{7}\,M_{\odot}$, which amounts to $\sim 80\%$ of the GMC mass converted into stars.
@@
5 Discussion
5.1 Wolf-Rayet stars and stellar metallicity
@@
Moreover, based on the C/O abundance ratio, Tapia et al. (2024) excluded the presence of WR stars in the system to avoid over-production of carbon, while this seems incompatible with the detection of WR spectral features (Rivera-Thorsen et al., 2024). From PARSEC data (Costa et al., 2025, see § 2), we found that the yields of both carbon and nitrogen have non-monotonic metallicity dependence. In Fig. 2, we also present the chemical composition of the stellar winds, where the C/O ratio can exceed $\log(\rm C/O)=0$ after $\sim 3\,\rm Myr$ for most metallicities in the plot except $Z=0.002$ and $Z=0.004$ (note that the inferred metallicity of the LyC cluster is in this range). As a result, the observed level of C enrichment by WR stars is reproduced by our simulations, as WR stars can exist in the cluster without depositing too much carbon due to the metallicity dependence of chemical yields. Similarly, Mollá & Terlevich (2012) calculated the cumulative element yields of stellar populations for different metallicity values. They found that the carbon abundance ($12+\log(\rm C/H)$, see their Fig. 9) in stellar winds is strongly enriched for $Z\geq 0.08$, but is actually slightly decreased for $Z\leq 0.004$. These results suggest that WR winds do not overproduce carbon at intermediate metallicity.
@@
Mollá & Terlevich (2012) also found that N/H is enhanced by $\sim 1$ dex due to stellar winds for all metallicities (while the enhancement in C/H only appears for $Z>0.2\,Z_{\odot}$). The oxygen abundance, however, can grow by $\sim 1$ dex for $Z=0.008$ and $Z=0.02$, but only slightly declines with time for $Z\leq 0.004$. This implies that a high N/O due to stellar winds is only possible for $Z\leq 0.004$. We note that this behavior is in agreement with our compilation based on Costa et al. (2025) in § 2.
@@
Interestingly, from a comprehensive compilation of the nitrogen-enriched star clusters and high-redshift galaxies (Ji et al., 2025, see their Table 2), the metallicities of these objects (including the Sunburst LyC cluster) are typically around $12+\log(\rm O/H)\sim 7.5-8$ (also see $z<0.5$ N-emitters from DESI DR1, Bhattacharya & Kobayashi, 2025). This could be a reflection of the intrinsic metallicity-dependent stellar evolution that affects the element yields of massive star winds (Mollá & Terlevich, 2012).
@@
We argue that this may be related to the yield during the WR phase of massive stars (and VMSs), which is metallicity-dependent (Maeder, 1992). In this phase, the stellar core undergoes helium burning (triple-alpha reaction), producing carbon and oxygen, while the shell/envelope is hydrogen-burning (via CNO cycle) and nitrogen-rich. At high metallicity, high mass loss rates due to strong stellar winds lead to stripping of material from first the envelope and then the core, explaining spectral evolution from WN to WC and WO stars. The envelope- and core- stripping phases combined leads to ejection of sizable amounts of nitrogen, carbon and oxygen (Maeder, 1992; Higgins et al., 2024). However, at low metallicity, WC and WO stars are rarer (Eldridge & Vink, 2006; Crowther, 2007), as the weaker stellar winds (Tramper et al., 2016) can only strip the nitrogen-rich envelope but hardly the core (Aguilera-Dena et al., 2022), leading to a very high N/O abundance ratio in the winds and a much reduced amount of carbon yield.
@@
In summary, we argue that the WR stars, as the late evolutionary phase of massive stars, and VMSs, should have formed in the Sunburst LyC cluster. Moreover, the SMC-like metallicity of nitrogen-enriched galaxies and clusters (Ji et al., 2025) can be understood from the evolutionary models of WR stars (Aguilera-Dena et al., 2022).
5.2 High-redshift nitrogen-enriched galaxies
@@
Recent advances in observations, especially with JWST, have uncovered that a small fraction (a few percent) of high-redshift ($z\sim 5$1111) galaxies exhibit strong nebular emission lines indicating an elevated gas-phase nitrogen abundance (e.g., Cameron et al., 2023; Bunker et al., 2023; Isobe et al., 2023; Navarro-Carrera et al., 2024; Schaerer et al., 2024; Marques-Chaves et al., 2024; Ji et al., 2024; Topping et al., 2025a). These “N-emitters” feature high N/O ratios, with some even reaching $\log(\rm N/O)\gtrsim 0.5$ (e.g., Isobe et al., 2023; Ji et al., 2024). Moreover, N-emitters can be associated with starbursts (Álvarez-Márquez et al., 2025).
@@
We speculate that the physical processes localized to newborn super star clusters as discovered in this work could be responsible for the spectral behavior of these high-redshift N-emitters. VMSs are theoretically expected to exist in the early Universe (Schaerer et al., 2025), and at least an excess of $50\,M_{\odot}$ massive stellar population has been observed (Cullen et al., 2025). Once starbursts happen in galaxies, the newly formed young stars may enrich nearby gas through stellar winds. The enriched gas may not be representative of the mean ISM properties of the whole galaxy, but instead only a small amount of high-pressure, dense photo-ionized gas in the proximity of highly massive, compact star clusters. This can explain why the nitrogen-enriched nebulae are compact in size (e.g., Topping et al., 2024, 2025a). We also note that several observations (Isobe et al., 2023; Senchyna et al., 2024; Ji et al., 2025) found chemical similarities between N-emitters and nitrogen-rich globular cluster stars, suggesting possibly an underlying connection between these two contexts.
@@
McClymont et al. (2025) attribute ISM nitrogen enhancement in galaxies to recent starbursts. The primary sources of nitrogen in their model are the AGB stars. Similarly, chemical evolution models by Rizzuti et al. (2025) found that differential winds can lead to the enrichment, while oxygen and $\alpha$ elements are produced by CCSNe. These different scenarios, as well as the VMS scenario presented here, may need further detailed studies to confirm in the context of high-redshift N-emitters. However, both AGB stars and CCSNe are typically much older than $4\,\rm Myr$, so we argue that they are unlikely to explain the nitrogen enrichment of the Sunburst Arc LyC cluster. The winds are also unlikely to be as dense as the observed density of $\gtrsim 10^{5}\,\rm cm^{-3}$ for some high-redshift N-emitters (Pascale et al., 2023; Ji et al., 2024; Yanagisawa et al., 2024).
@@
Finally, from our simulation, star clusters here can have high central densities ($\sim 10^{5}\,M_{\odot}\,{\rm pc}^{-3}$), so another potential source in the nitrogen mass budget may be supermassive stars (Charbonnel et al., 2023; Nandal et al., 2025; Gieles et al., 2025) as the product of runaway stellar mergers and the transition to intermediate-mass black holes (IMBHs; e.g., Portegies Zwart & McMillan, 2002; Gürkan et al., 2004; Kremer et al., 2020; Shi et al., 2021; Fujii et al., 2024). However, the relative importance of this channel may require a closer look given that VMSs alone can already explain the N/O anomaly. Particularly for the Sunburst Arc, the VMSs and their mass segregation are detected (Meštrić et al., 2023), while a supermassive star should feature a central bright point source, so we argue that supermassive stars may not be the explanation for this object.
@@
5.3 Nitrogen-enriched AGNs
Observations also found high-redshift, nitrogen-enriched broad-line AGNs (Isobe et al., 2025), which can also be related to this scenario, that the broad-line region (BLR) is polluted by “field stars” in the nuclear star cluster after a starburst. However, we caution that the source of this enrichment can also be related to star formation inside AGN disks, which undergoes a completely different evolution track compared with field stars (Cantiello et al., 2021; Jermyn et al., 2022; Ali-Dib & Lin, 2023). For example, Huang et al. (2023) found that apart from the nitrogen enrichment in the BLR, there is also significant C+N+O enrichment. This differs from the globular cluster situation where C+N+O is roughly uniform among multiple stellar populations (which is the case for narrow-line region; Huang et al., 2023), and hence implies a different chemical feedback process (Fryer et al., 2025).
Fig. 13 shows the abundance of different elements of the high-pressure, nitrogen-enriched gas from our fiducial star cluster simulations (§ 4.1). The abundances of C and O exhibit unremarkable deviations from the natal values (at most $\sim 0.2\,$ dex). Apart from the significant N enrichment, it is interesting to see a significant elevation in He abundance (0.10.10.20.2\,dex). However, the sum C+N+O is very similar to the natal value, which is unlike the case of BLR enrichment in Huang et al. (2023). VMSs may or may not play a role in the nitrogen enrichment of AGNs.
6 Conclusion
By semi-analytic calculations (§ 2) and running magnetohydrodynamical simulations with radiative feedback (§ 3, 4), we have presented evidence that nitrogen enrichment of the high-pressure gas in the Sunburst LyC cluster is the combined result of both the formation of massive compact star cluster and chemical feedback from very massive stars. The enrichment source can be a hybrid of both Vink (2023, VMSs only) and Kobayashi & Ferrara (2024, WR stars with non-VMS origin). But a key difference is that the effect we study operates on the star cluster scales rather than galaxy scales, given the observed stellar wind signatures of VMSs and/or WR stars (Meštrić et al., 2023; Rivera-Thorsen et al., 2024). We summarize below the key physical processes of cluster formation and feedback:
@@ 6 Conclusion
    Super star cluster formation from dense GMC collapse. For an initial GMC with a mean surface density as high as $\sim 10^{3}$$10^{4}\,M_{\odot}\,{\rm pc}^{-2}$, a star cluster may form at a very high star formation efficiency of $\sim 30\%$, leading to a super star cluster (consisting of a central cluster and many satellites) of $\sim 10^{7}\,M_{\odot}$ that matches the inferred stellar mass of the Sunburst LyC cluster (Vanzella et al., 2022; Pascale et al., 2023; Rivera-Thorsen et al., 2024). If the initial GMC is more compact, hierarchical mergers build up an even more mass-dominating central cluster (Shi et al., 2024). The compact star cluster can confine the gas in its cluster.
@@
    Gas sinking under the pull of gravity. Within the first global free-fall time of the GMC, cold natal gas is pulled in by gravity toward the cluster center. The density of this gas dramatically increases ($\gtrsim 10^{5}\,\rm cm^{-3}$) through this nearly isothermal gravitational collapse. We identify this as a new dynamical reason to account for the observation that nitrogen-enriched gas is denser (and more pressurized) than typical warm ionized gas in the ISM (e.g., Pascale et al., 2023; Ji et al., 2024; Yanagisawa et al., 2024).
@@
    Chemical yields of VMS winds. Stellar winds from massive stars, especially from VMSs, are the source of pollution that enriches the dense gas pulled to the cluster center. Particularly for the Sunburst LyC cluster, the VMS population at $Z=0.2\,Z_{\odot}$ enrich the natal gas that has fallen to the center, such that $\log(\rm N/O)\gtrsim-0.5$ while carbon is not over-abundant. This matches what is measured from the observed high-pressure nebula (Pascale et al., 2023). We also predict that the dense gas exhibits 0.10.10.20.2\,dex of helium abundance elevation.
@@
    Stellar irradiation. Simultaneously with chemical enrichment, the dense gas becomes photo-heated and ionized, and is subject to strong radiation pressure, while kinetic feedback from winds and SNe is secondary (§ 4.4). Radiation pressure therefore is an important contributor to the high pressure of the nitrogen-enriched gas, while most of the remaining natal gas at larger radii has a low pressure and is only slightly enriched (Pascale et al., 2023) (i.e., stratification of nebular physical and chemical properties; Pascale et al., 2023; Ji et al., 2024; Topping et al., 2025b). Soon after enrichment of the innermost dense gas, which amounts to $\sim 10^{4}\,M_{\odot}$, strong irradiation efficiently ejects it out of the cluster potential, at around $\sim 4\,\rm Myr$.
@@
Massive compact star cluster formation requires natal GMCs with a high surface density. Our fiducial simulations suggest that GMCs weighing $3\times 10^{7}\,M_{\odot}$ and initially of a size $\sim 70\,\rm pc$ (mean surface density $\Sigma\sim 2\times 10^{3}\,M_{\odot}\,{\rm pc}^{-2}$) can collapse to give the desired cluster mass and radius as observed for the Sunburst LyC cluster, while simulations with lower initial surface density values do not due to a lower star formation efficiency (Grudić et al., 2018; Chevance et al., 2023). This requirement of high surface density agrees with other simulation results that focus on the escape of LyC photons from massive compact star clusters (Menon et al., 2025). Moreover, what forms out of the natal GMC is not a single star cluster, but instead a central massive cluster surrounded by many satellite sub-clusters (Fig. 7).
@@
Very Massive Stars are required to explain the rapid, heavy enrichment of nitrogen in the Sunburst LyC cluster at $\sim 2$$4\,\rm Myr$. Though OB stars might lead to similar nitrogen enrichment, the time required to reach the observed high N/O value would be too long compared to observations, and this scenario would also require the absence of WR stars (Tapia et al., 2024). This is in tension with spectroscopy (Meštrić et al., 2023; Rivera-Thorsen et al., 2024).
@@
Wolf-Rayet stars, specifically WN-type stars, as a late and helium-burning stage of massive stars, and VMSs, should either be present or have formed before in the LyC cluster (Rivera-Thorsen et al., 2024). At the inferred metallicity values of nitrogen-enriched galaxies and super star clusters ($12+\log(\rm O/H)\sim 7.5-8$, or $Z\lesssim 0.2\,Z_{\odot}$, see Ji et al., 2025), stellar winds efficiently shed the N-rich envelope of WN stars while carbon and oxygen are trapped in the stellar core (Aguilera-Dena et al., 2022), leading to a high N/O but slightly elevated C/O in the wind ejecta. Both our yield model (§ 2) and Mollá & Terlevich (2012) find this metallicity dependence in terms of N/O and C/O ratios.
@@
We suggest that in high-redshift nitrogen-enriched galaxies like GN-z11 and GS_3073, N/O elevation may represent a temporary phase during the bursty formation of dense stellar systems. Compared to the scenario of differential winds (Rizzuti et al., 2025; McClymont et al., 2025), this is associated with gas trapped in the proximity of newborn super star clusters (galaxies with N/O anomaly do have high surface density $\Sigma_{\star}\gtrsim 10^{2.5}\,M_{\odot}\,{\rm pc}^{-2}$, see Ji et al., 2025). The VMS population provides the extra nitrogen, or more generally speaking, WR stars at sub-solar metallicity (Fig. 2). This enrichment, occurring in dense gas that is pressurized and photoionized by young stars, can explain the observed nitrogen nebular emission lines. By contrast, the majority of the gas is low-density, low-pressure gas, which is not enriched (Ji et al., 2024).
@@
One prediction of this scenario is a significant, 0.10.10.20.2\,dex elevation of the He/H ratio in the high-pressure gas that would accompany the N/O anomaly (see Fig. 5), which can be measurable and explain the positive He/H-N/O correlation (Yanagisawa et al., 2024). This enrichment is also due to winds from massive stars including the VMSs, when the envelope of these stars becomes helium-elevated in their WR stages. Similar to the case of the N/O abundance ratio, the majority of the gas, being low-pressure, is more dominated by the natal gas components and is hardly helium-enhanced.

### v2

@@ IV.4 VMS feedback tests
Therefore, the only remaining strong feedback mechanism has to do with radiation. In the third test, we disable the radiative feedback from all stars (the VMS population alone elevates the overall L/M ratio, but only by a factor of $\sim 1.5$, see § III). With the third row in Fig. 13, we see that radiative feedback indeed makes a big difference. The SFE is much higher as the cumulative stellar mass reaches $\sim 2.5\times 10^{7}\,M_{\odot}$, which amounts to $\sim 80\%$ of the GMC mass converted into stars.
@@
V Discussion
V.1 Wolf-Rayet stars and stellar metallicity
@@
Moreover, based on the C/O abundance ratio, TapiaBekkiGroves_2024MNRAS.534.2086T excluded the presence of WR stars in the system to avoid over-production of carbon, while this seems incompatible with the detection of WR spectral features (Rivera-ThorsenChisholmWelch_2024A&A...690A.269R). From PARSEC data (CostaShepherdBressan_2025A&A...694A.193C, see § II), we found that the yields of both carbon and nitrogen have non-monotonic metallicity dependence. In Fig. 3, we also present the chemical composition of the stellar winds, where the C/O ratio can exceed $\log(\rm C/O)=0$ after $\sim 3\,\rm Myr$ for most metallicities in the plot except $Z=0.002$ and $Z=0.004$ (note that the inferred metallicity of the LyC cluster is in this range). As a result, the observed level of C enrichment by WR stars is reproduced by our simulations, as WR stars can exist in the cluster without depositing too much carbon due to the metallicity dependence of chemical yields. Similarly, MollaTerlevich_2012MNRAS.425.1696M calculated the cumulative element yields of stellar populations for different metallicity values. They found that the carbon abundance ($12+\log(\rm C/H)$, see their Fig. 9) in stellar winds is strongly enriched for $Z\geq 0.08$, but is actually slightly decreased for $Z\leq 0.004$. These results suggest that WR winds do not overproduce carbon at intermediate metallicity.
@@
MollaTerlevich_2012MNRAS.425.1696M also found that N/H is enhanced by $\sim 1$ dex due to stellar winds for all metallicities (while the enhancement in C/H only appears for $Z>0.2\,Z_{\odot}$). The oxygen abundance, however, can grow by $\sim 1$ dex for $Z=0.008$ and $Z=0.02$, but only slightly declines with time for $Z\leq 0.004$. This implies that a high N/O due to stellar winds is only possible for $Z\leq 0.004$. We note that this behavior is in agreement with our compilation based on CostaShepherdBressan_2025A&A...694A.193C in § II.
@@
Interestingly, from a comprehensive compilation of the nitrogen-enriched star clusters and high-redshift galaxies (JiBelokurovMaiolino_2025MNRAS.tmp.1982J, see their Table 2), the metallicities of these objects (including the Sunburst LyC cluster) are typically around $12+\log(\rm O/H)\sim 7.5-8$ (also see $z<0.5$ N-emitters from DESI DR1, BhattacharyaKobayashi_2025arXiv250811998B). textcolorblackMore recent data of N-emitters from MorelSchaererMarques-Chaves_2025arXiv251120484M revealed a wider range of $12+\log(\rm O/H)\sim 7-8.5$, yet still sub-solar. Moreover, the C/O abundance of these N-emitters (including the Sunburst Arc, see Fig. 2) aligns with the local relation. This could be a reflection of the intrinsic metallicity-dependent stellar evolution that affects the element yields of massive star winds (MollaTerlevich_2012MNRAS.425.1696M).
@@
We suggest that the trend arises from metallicity-dependent yields during the WR phase of massive stars and VMSs (Maeder_1992A&A...264..105M). During this phase, helium burning in the core (via the triple-alpha reaction) produces carbon and oxygen, while hydrogen burning in the shell (through the CNO cycle) enriches nitrogen. At high metallicity, strong winds strip first the envelope and then the core, driving the spectral evolution from WN to WC/WO stars and ejecting substantial nitrogen, carbon, and oxygen (Maeder_1992A&A...264..105M; HigginsVinkHirschi_2024MNRAS.533.1095H). At low metallicity, weaker winds (TramperSanadeKoter_2016ApJ...833..133T) remove mainly the nitrogen-rich envelope but rarely the core (Aguilera-DenaLangerAntoniadis_2022A&A...661A..60A), making WC/WO stars rarer (EldridgeVink_2006A&A...452..295E; Crowther_2007ARA&A..45..177C) and yielding winds with high N/O and little carbon.
@@
In summary, WR stars and VMSs likely formed in the Sunburst LyC cluster. The SMC-like metallicities observed in nitrogen-enriched galaxies and clusters (JiBelokurovMaiolino_2025MNRAS.tmp.1982J) are consistent with WR evolutionary models (Aguilera-DenaLangerAntoniadis_2022A&A...661A..60A).
V.2 High-redshift nitrogen-enriched galaxies
@@
Recent advances in observations, especially with JWST, have uncovered that a small fraction (a few percent) of high-redshift ($z\sim 5$1111) galaxies exhibit strong nebular emission lines indicating an elevated gas-phase nitrogen abundance (e.g., CameronKatzRey_2023MNRAS.523.3516C; BunkerSaxenaCameron_2023A&A...677A..88B; IsobeOuchiTominaga_2023ApJ...959..100I; Navarro-CarreraCaputiIani_2025ApJ...993..194N; SchaererMarques-ChavesXiao_2024A&A...687L..11S; Marques-ChavesSchaererKuruvanthodi_2024A&A...681A..30M; JiUblerMaiolino_2024MNRAS.535..881J; ToppingStarkSenchyna_2025ApJ...980..225T). These “N-emitters” feature high N/O ratios, with some even reaching $\log(\rm N/O)\gtrsim 0.5$ (e.g., IsobeOuchiTominaga_2023ApJ...959..100I; JiUblerMaiolino_2024MNRAS.535..881J). Moreover, N-emitters can be associated with starbursts (Alvarez-MarquezCrespoGomezColina_2025A&A...695A.250A).
@@
We speculate that the physical processes localized to newborn super star clusters as discovered in this work could be responsible for the spectral behavior of these high-redshift N-emitters. VMSs are theoretically expected to exist in the early Universe (SchaererGuibertMarques-Chaves_2025A&A...693A.271S), and at least an excess of $50\,M_{\odot}$ massive stellar population has been observed (CullenCarnallScholte_2025MNRAS.540.2176C). Once starbursts happen in galaxies, the newly formed young stars may enrich nearby gas through stellar winds. The enriched gas may not be representative of the mean ISM properties of the whole galaxy, but instead only a small amount of high-pressure, dense photo-ionized gas in the proximity of highly massive, compact star clusters. This can explain why the nitrogen-enriched nebulae are compact in size (e.g., ToppingStarkSenchyna_2024MNRAS.529.3301T; ToppingStarkSenchyna_2025ApJ...980..225T). We also note that several observations (IsobeOuchiTominaga_2023ApJ...959..100I; SenchynaPlatStark_2024ApJ...966...92S; JiBelokurovMaiolino_2025MNRAS.tmp.1982J) found chemical similarities between N-emitters and nitrogen-rich globular cluster stars, suggesting possibly an underlying connection between these two contexts.
@@
McClymontTacchellaSmith_2025arXiv250708787M attribute ISM nitrogen enhancement in galaxies to recent starbursts and AGB stars. Similarly, chemical evolution models by RizzutiMatteucciMolaro_2025A&A...697A..96R found that differential winds can lead to the enrichment, while oxygen and $\alpha$ elements are produced by CCSNe. These different scenarios, as well as the VMS scenario presented here, may need further detailed studies to confirm in the context of high-redshift N-emitters. However, both AGB stars and CCSNe are typically much older than $4\,\rm Myr$, so we argue that they are unlikely to explain the nitrogen enrichment of the Sunburst Arc LyC cluster. The winds are also unlikely to be as dense as the observed density of $\gtrsim 10^{5}\,\rm cm^{-3}$ for some high-redshift N-emitters (PascaleDaiMcKee_2023ApJ...957...77P; JiUblerMaiolino_2024MNRAS.535..881J; YanagisawaOuchiWatanabe_2024ApJ...974..266Y).
@@
Finally, from our simulation, star clusters here can have high central densities ($\sim 10^{5}\,M_{\odot}\,{\rm pc}^{-3}$), so another potential source in the nitrogen mass budget may be supermassive stars (CharbonnelSchaererPrantzos_2023A&A...673L...7C; NandalWhalenLatif_2025ApJ...994L..11N; GielesPadoanCharbonnel_2025MNRAS.544..483G) as the product of runaway stellar mergers and the transition to intermediate-mass black holes (IMBHs; e.g., PortegiesZwartMcMillan_2002ApJ...576..899P; GurkanFreitagRasio_2004ApJ...604..632G; KremerSperaBecker_2020ApJ...903...45K; ShiGrudicHopkins_2021MNRAS.505.2753S; FujiiWangTanikawa_2024Sci...384.1488F). However, the relative importance of this channel may require a closer look given that VMSs alone can already explain the N/O anomaly. Particularly for the Sunburst Arc, the VMSs and their mass segregation are detected (MestricVanzellaUpadhyaya_2023A&A...673A..50M), while a supermassive star should feature a central bright point source, so we argue that supermassive stars may not be the explanation for this object.
@@
V.3 Nitrogen-enriched AGNs
Observations also found high-redshift, nitrogen-enriched broad-line AGNs (IsobeMaiolinoD'Eugenio_2025MNRAS.541L..71I), which can also be related to this scenario, that the broad-line region (BLR) is polluted by “field stars” in the nuclear star cluster after a starburst. However, we caution that the source of this enrichment can also be related to star formation inside AGN disks, which undergoes a completely different evolution track compared with field stars (CantielloJermynLin_2021ApJ...910...94C; JermynDittmannMcKernan_2022ApJ...929..133J; Ali-DibLin_2023MNRAS.526.5824A). For example, HuangLinShields_2023MNRAS.525.5702H found that apart from the nitrogen enrichment in the BLR, there is also significant C+N+O enrichment. This differs from the globular cluster situation where C+N+O is roughly uniform among multiple stellar populations (which is the case for narrow-line region; HuangLinShields_2023MNRAS.525.5702H), and hence implies a different chemical feedback process (FryerHuangAli-Dib_2025MNRAS.537.1556F).
Fig. 14 shows the abundance of different elements of the high-pressure, nitrogen-enriched gas from our fiducial star cluster simulations (§ IV.1). The abundances of C and O exhibit unremarkable deviations from the natal values (at most $\sim 0.2\,$ dex). Apart from the significant N enrichment, it is interesting to see a significant elevation in He abundance (0.10.10.20.2\,dex). However, the sum C+N+O is very similar to the natal value, which is unlike the case of BLR enrichment in HuangLinShields_2023MNRAS.525.5702H. VMSs may or may not play a role in the nitrogen enrichment of AGNs.
VI Conclusion
By semi-analytic calculations (§ II) and running magnetohydrodynamical simulations with radiative feedback (§ III, IV), we have presented evidence that nitrogen enrichment of the high-pressure gas in the Sunburst LyC cluster is the combined result of both the formation of massive compact star cluster and chemical feedback from very massive stars. The enrichment source can be a hybrid of both Vink_2023A&A...679L...9V and KobayashiFerrara_2024ApJ...962L...6K. But a key difference is that the effect we study operates on the star cluster scales rather than galaxy scales, given the observed stellar wind signatures of VMSs and/or WR stars (MestricVanzellaUpadhyaya_2023A&A...673A..50M; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R). We summarize below the key physical processes of cluster formation and feedback:
@@ VI Conclusion
    Super star cluster formation from dense GMC collapse. For an initial GMC with a mean surface density as high as $\sim 10^{3}$$10^{4}\,M_{\odot}\,{\rm pc}^{-2}$, a star cluster may form at a very high star formation efficiency of $\sim 30\%$, leading to a super star cluster (consisting of a central cluster and many satellites) of $\sim 10^{7}\,M_{\odot}$ that matches the inferred stellar mass of the Sunburst LyC cluster (VanzellaCastellanoBergamini_2022A&A...659A...2V; PascaleDaiMcKee_2023ApJ...957...77P; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R). If the initial GMC is more compact, hierarchical mergers build up an even more mass-dominating central cluster (ShiKremerHopkins_2024ApJ...969L..31S). The compact star cluster can confine the gas in its cluster.
@@
    Gas sinking under the pull of gravity. Within the first global free-fall time of the GMC, cold natal gas is pulled in by gravity toward the cluster center. The density of this gas dramatically increases ($\gtrsim 10^{5}\,\rm cm^{-3}$) through this nearly isothermal gravitational collapse. We identify this as a new dynamical reason to account for the observation that nitrogen-enriched gas is denser (and more pressurized) than typical warm ionized gas in the ISM (e.g., PascaleDaiMcKee_2023ApJ...957...77P; JiUblerMaiolino_2024MNRAS.535..881J; YanagisawaOuchiWatanabe_2024ApJ...974..266Y).
@@
    Chemical yields of VMS winds. Stellar winds from massive stars, especially from VMSs, are the source of pollution that enriches the dense gas pulled to the cluster center. Particularly for the Sunburst LyC cluster, the VMS population at $Z=0.2\,Z_{\odot}$ enrich the natal gas that has fallen to the center, such that $\log(\rm N/O)\gtrsim-0.5$ while carbon is not over-abundant. This matches what is measured from the observed high-pressure nebula (PascaleDaiMcKee_2023ApJ...957...77P). We also predict that the dense gas exhibits 0.10.10.20.2\,dex of helium abundance elevation.
@@
    Stellar irradiation. Simultaneously with chemical enrichment, the dense gas becomes photo-heated and ionized, and is subject to strong radiation pressure, while kinetic feedback from winds and SNe is secondary (§ IV.4). Radiation pressure therefore is an important contributor to the high pressure of the nitrogen-enriched gas, while most of the remaining natal gas at larger radii has a low pressure and is only slightly enriched (PascaleDaiMcKee_2023ApJ...957...77P) (i.e., stratification of nebular physical and chemical properties; PascaleDaiMcKee_2023ApJ...957...77P; JiUblerMaiolino_2024MNRAS.535..881J; ToppingSandersShapley_2025MNRAS.541.1707T). Soon after enrichment of the innermost dense gas, which amounts to $\sim 10^{4}\,M_{\odot}$, strong irradiation efficiently ejects it out of the cluster potential, at around $\sim 4\,\rm Myr$.
@@
Massive compact star cluster formation requires natal GMCs with a high surface density. Our fiducial simulations suggest that GMCs weighing $3\times 10^{7}\,M_{\odot}$ and initially of a size $\sim 70\,\rm pc$ (mean surface density $\Sigma\sim 2\times 10^{3}\,M_{\odot}\,{\rm pc}^{-2}$) can collapse to give the desired cluster mass and radius as observed for the Sunburst LyC cluster, while simulations with lower initial surface density values do not due to a lower star formation efficiency (GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G; ChevanceKrumholzMcLeod_2023ASPC..534....1C). This requirement of high surface density agrees with other simulation results that focus on the escape of LyC photons from massive compact star clusters (MenonBurkhartSomerville_2025ApJ...987...12M). Moreover, what forms out of the natal GMC is not a single star cluster, but instead a central massive cluster surrounded by many satellite sub-clusters (Fig. 8).
@@
Very Massive Stars are required to explain the rapid, heavy enrichment of nitrogen in the Sunburst LyC cluster at $\sim 2$$4\,\rm Myr$. Though OB stars might lead to similar nitrogen enrichment, the time required to reach the observed high N/O value would be too long compared to observations, and this scenario would also require the absence of WR stars (TapiaBekkiGroves_2024MNRAS.534.2086T). This is in tension with spectroscopy (MestricVanzellaUpadhyaya_2023A&A...673A..50M; Rivera-ThorsenChisholmWelch_2024A&A...690A.269R).
@@
Wolf-Rayet stars, specifically WN-type stars, as a late and helium-burning stage of massive stars, and VMSs, should either be present or have formed before in the LyC cluster (Rivera-ThorsenChisholmWelch_2024A&A...690A.269R). At the inferred metallicity values of nitrogen-enriched galaxies and super star clusters ($12+\log(\rm O/H)\sim 7.5-8$, or $Z\lesssim 0.2\,Z_{\odot}$, see JiBelokurovMaiolino_2025MNRAS.tmp.1982J), stellar winds efficiently shed the N-rich envelope of WN stars while carbon and oxygen are trapped in the stellar core (Aguilera-DenaLangerAntoniadis_2022A&A...661A..60A), leading to a high N/O but slightly elevated C/O in the wind ejecta. Both our yield model (§ II) and MollaTerlevich_2012MNRAS.425.1696M find this metallicity dependence in terms of N/O and C/O ratios.
@@
We suggest that in high-redshift nitrogen-enriched galaxies like GN-z11 and GS_3073, N/O elevation may represent a temporary phase during the bursty formation of dense stellar systems. Compared to the scenario of differential winds (RizzutiMatteucciMolaro_2025A&A...697A..96R; McClymontTacchellaSmith_2025arXiv250708787M), this is associated with gas trapped in the proximity of newborn super star clusters (galaxies with N/O anomaly do have high surface density $\Sigma_{\star}\gtrsim 10^{2.5}\,M_{\odot}\,{\rm pc}^{-2}$, see JiBelokurovMaiolino_2025MNRAS.tmp.1982J). The VMS population provides the extra nitrogen, or more generally speaking, WR stars at sub-solar metallicity (Fig. 3). This enrichment, occurring in dense gas that is pressurized and photoionized by young stars, can explain the observed nitrogen nebular emission lines. By contrast, the majority of the gas is low-density, low-pressure gas, which is not enriched (JiUblerMaiolino_2024MNRAS.535..881J).
@@
One prediction of this scenario is a significant, 0.10.10.20.2\,dex elevation of the He/H ratio in the high-pressure gas that would accompany the N/O anomaly (see Fig. 6), which can be measurable and explain the positive He/H-N/O correlation (YanagisawaOuchiWatanabe_2024ApJ...974..266Y). This enrichment is also due to winds from massive stars including the VMSs, when the envelope of these stars becomes helium-elevated in their WR stages. Similar to the case of the N/O abundance ratio, the majority of the gas, being low-pressure, is more dominated by the natal gas components and is hardly helium-enhanced.